OpenStudy (anonymous):

I GIVE GOLD MEDAL. generate the first 5 terms of the sequence of f(1)=1 and f(2)=2, f(n)=f(n-1)+f(n-2), for n > 2 a. 1, 2, 3, 4, 5 b. 1, 2, 3, 5, 7 c. 1, 2, 3, 5, 8 d. 1, 3, 5, 7, 9

OpenStudy (anonymous):

Watch

OpenStudy (anonymous):

Plug in n = 3, f(3) = f(3-1) + f(3 - 2) = f(2) + f(1) = 2 + 1 = 3

OpenStudy (anonymous):

Now plug in n = 4, you get f(f) = f(4-1) + f(4-2) = f(3) + f(2) = 3 + 2 = 5

OpenStudy (anonymous):

So f(4) = 5